# Realistic binary neutron star initial data with Elliptica

### v1

@@ 1 Introduction
Simulation of compact binary system in NR often involves two steps. The first step is to find the solution of Einstein-Euler PDE on a hypersurface of the spacetime manifold, namely, constructing constraint satisfying and self consistent Initial Data (ID) that present the binary system of interest at some time. ID codes such as, COCAL [22, 23, 24], Elliptica [25], FUKA [26], LORENE [27, 28, 29, 30, 31], NRPyElliptic [32], SGRID [33, 34, 35], SpECTRE’s elliptic solver [36, 37], Spells [38, 39, 40, 41], TwoPunctures [42, 43], are developed and utilized for this step.
The next step is to use the ID and simulate the system’s evolution over time; hence we can find the solution over the spacetime of interest. Dynamical evolution codes, for instance, BAM [44, 45, 46], BAMPS [47, 48, 49], Dendro-GR [50], Einstein Toolkit [51, 52], ExaHyPE [53] GR-Athena++ [54, 55, 56] GRaM-X [57, 58], GRChombo [59, 60], Nmesh [61], NRPy+ [62], SACRA-MPI [63], Simflowny [64], SpEC [65, 66], SpECTRE [67, 68], and SPHINCS_BSSN [69], among others, are made and employed for this step.
@@
The remainder of paper is organized as follows. In section 2 we explain the mathematical background and formalism we use for ID of BNS systems, in particular, Einstein-Euler equations and EOS in Elliptica. Section 3 details the underling algorithms and numerical techniques for construction of BNS’s ID. We present the implementation of tabulated EOS by spline interpolant means and diagnostics in Elliptica. Additionally, we explain the iterative procedure for finding the physical and constraint satisfying ID for BNS systems. In section 4, we present various convergence tests and comparison against post Newtonian answers to showcase the proof of concept for the new version the code. In section 5, we discuss the possible improvements and future work.
Throughout the article, we use geometric units where $G=1$ is the constant of gravity, $c=1$ is the speed of light, and solar mass $M_{\odot}=1$.
@@ 2.1 Einstein-Euler equations
Here $\psi$ is the conformal factor, $\bar{\gamma}_{ij}$ the conformal 3-metric, $A^{ij}$ the traceless part of $K^{ij}$. Moreover, since we are using the Extended Conformal Thin Sandwich (XCTS) formalism [74, 73], $A^{ij}$ is decomposed as
@@
here, $\Omega^{j}_{\mathrm{NS}}$ is a free parameter to set the spin level, see sec. 4.3, and $x^{k}_{c}$ denotes the NS coordinate. The irrotational part is shown by the potential $\phi(x,y,z)$ and obeys the following equations. [35]
@@ 3.1 Coordinate system
Figure 1: Grid patches for a BNS system. Left: shown is the $x-y$ plane of the computational grid that is covered by different patches. Right: a zoomed-in view of the $x-y$ plane focusing on the NS regions. A Cartesian box is used around each NS center to avoid coordinate singularities. By using various $\sigma(X,Y)$ values, cubed spherical patches can adapt to different shapes, effectively capturing the shape of the NS surface. This allows for the treatment of matter and vacuum in separate patches.
Fig. (1) illustrates the patches covering the $x-y$ plane of the computational grid used in this work.
@@ 3.2 Elliptic solver
Given an elliptic PDE, we linearize the equation to use the Newton-Raphson method (see [25] for a complete description). Thus, the original elliptic PDE becomes a matrix equation, like $Ax=B$, to be solved. Now the challenge is to solve this matrix equation efficiently. To this aim, we use the Schur domain decomposition method [80, 25]. At the core of this method there is a divide and conquer strategy; in this strategy, first the coupled equations are solved and then the system decomposes into independent (uncoupled) subsystems, amenable for parallel solving. The Schur domain decomposition allows to solve the full system using shared memory multiprocessing in which each CPU core is assigned to solve a matrix equations at each patch independently.
The Schur domain decomposition method arranges the Jacobian matrix in Newton-Raphson method into two parts. The first part is associated with all uncoupled equations and the second part comprises of the coupled equations. These couplings between equations occur due to the inter-BC at the interfaces of adjacent patches [25]. As such, the system of matrix equations can be seen as two equations with two unknowns as follows
@@
Therefore, we first solve for $\tilde{w}$ and then we use it to solve for $\tilde{v}$. Accordingly, in Elliptica, we solve eq. 35 like:
@@
As we can see eq (36) only involves the unknown $\tilde{w}$ and it can be summarized as
@@ 3.4 Tabulated equations of state
A number of slight modifications may be made to the EOS in pre-processing in order to make it more amenable to interpolation. These modifications are necessary for both the physical consistency of the EOS and to facilitate the convergence of the solution.
The first modification involves adjusting specific enthalpy at the NS surface for the EOS tables in which the specific enthalpy falls below $1$. To this aim, we scale all $\rho_{0}$ points by the multiplicative constant $\eta=h_{\mathrm{min}}$ corresponding to the lowest value of $h$ in the table. Accordingly, $h$ points are scaled by the value $\eta^{-1}$, i.e:
@@
This alteration can be carried through to dynamical evolution codes by e.g. using the same EOS table or by scaling the rest-mass density (or baryon number density) by the same factor.
In addition, several other features of tabular EOS may pose problems for both interpolation and the use of specific enthalpy as the independent thermodynamic variable. Among these problems is the presence of a region near the surface of the NS where both $\frac{de}{dh}$ and $\frac{d\rho_{0}}{dh}$ diverge. This region is common to many EOS with a ‘crust’, including both tabular EOS such as SFHo [84] and piecewise polytropics [85]. While this region is unavoidable if we wish to accurately represent the equation of state via specific enthalpy, cf. [86, 87], it also limits the accuracy of the solution within the NS (see fig. 3).
@@
Figure 2: Comparison of changes with respect to the specific enthalpy ($h$) in total energy density ($e$) and the rest mass density ($\rho_{0}$) near the NS surface for three EOS: K96 (single polytrope), SLy (piecewise polytrope), and SFHo (table). A lower specific enthalpy corresponds to a point closer to the NS surface, with $h=1$ defining the surface itself.
Fig. 2 shows the resultant tabulated SFHo EOS after re-griding using a low-degree Hermite interpolant with comparison to EOS of K96 and SLy. The K96 EOS is a single polytrope, corresponding to $K=96.7$ and $\Gamma=2$ [88]. SLy is an approximate piecewise polytrope of tabulated SLy4, and therefore quantities such as $\frac{de}{dh}$ are not continuous in certain regions corresponding to transitions between piecewise segments [89]. SFHo, here restricted to the zero-temperature regime, shows similar discontinuities [84].
Examining the functions relevant to the ID generation, e.g. $\rho_{0}$ and $e$, we see that the single-polytrope model is considerably simpler and more amenable to numerical solution than the models with discontinuities. The difference is typically relevant to the surface of the NS (where $h=1$), as shown in the nearly vertical profile of $\frac{d\rho_{0}}{dh}$ and $\frac{de}{dh}$ in fig. (2). The sharp changes in $\frac{d\rho_{0}}{dh}$ and $\frac{de}{dh}$ with the piecewise polytropic and tabular EOS are not perfectly captured by the grid, and thus they limit the convergence of the constraint violations (see sec. 4.1).
@@ 3.5 Initial data construction
Step 1: We solve the elliptic equations iteratively in this specific order: first eq. (25) for the matter field, and then eqs. (17), (18), and (19) for the metric fields. This approach has been found to enhance the solution’s stability, particularly when solving the matter field first. The order of solving the metric fields themselves, however, appears to be less critical. Moreover, during each iteration step, we focus on solving a single elliptic equation while keeping the other fields fixed. In essence, the fixed fields act as source terms influencing the equation being solved. Finally, within the Newton-Raphson iterative method, we perform only one update step per field and then incorporating the newly solved value into the source terms for the next equation. In the following $\Xi$ denotes a field from the set $\{\phi,\psi,\alpha\psi,B^{i}\}$.
@@
Step 3: After updating all $\Xi$ fields, we see the baryon mass of each NS deviates, often by a few percents, from the target value. Additionally, for spinning NS, since at Step 0 we used TOV solution and then we added the spin vector to the NS, it is not surprising to observe the baryon mass is different from the prescribed target value. Moreover, the starting resolution is often coarse and again the baryon mass may change as we go to a higher resolution. If we do not account for these changes, the baryon mass deviates even further at later iterations and may lead to a code crash. To adjust the baryon mass, we note that $\rho_{0}=\rho_{0}(h)$, and as shown in eq. (31) the specific enthalpy depends on a constant $C$, i.e., $h=h(C)$, which implies $\rho_{0}=\rho_{0}(C)$. Therefore by using a root finder, we find the value of $C$ in eq. (40) such that the baryon mass remains the same as the target value.
@@
Step 6: Finding the orbital angular velocity, $\Omega^{z}_{\mathrm{BNS}}$, for a quasi-circular orbit demands full solution of general relativity. Since we stat with a Newtonian approximation for this value we need to refine it. Following [90], so called the force balance method, we first compute $\partial_{i}\ln\Gamma$ at each NS center, where $\Gamma$ computed as
@@
Step 7: At this point the specific enthalpy profile and hence NS surface are changed. We need to find the new location of NS surface for adjusting cubed spherical patches that are covering the NS – so we can separate matter and vacuum into different patches. Since some parts of the NS may need to extend to the patches that are currently covering vacuum, we extrapolate $h$ into these patches so the root finder can find where $h=1$. To this aim we use the following formula to extrapolate specific enthalpy to the vacuum
@@
here, $r$ denotes the coordinate distance from the NS center, and $c_{0}$ is a constant (typically $0.01$). We find the values of $a$, and $b$ by demanding $C^{1}$ continuity across the NS surface. Additionally, we use the same eq. 53 to extrapolate $\phi$ field outside the NS. This step is required when the NS surface is expanded and we want to interpolate $\phi$ from the current grid to a new grid.
@@
Step 11: We monitor the Hamiltonian and momentum constraints, eqs (55) and (56). We restart from Step 1 unless the constraints reach their truncation error and are level off. In this case, when the constraints reach a plateau, we stop the iterative process if there is no higher resolution demanded; otherwise we go to the next resolution and start from Step 1.
@@ 4.1 Convergence test
For the convergence test, we generate ID for two symmetric BNS systems: one with polytropic K96 EOS and the other with tabulated SFHo EOS. In these systems, the NS have baryon mass $1.4$ with no spin, and their separation is $50$ units.
@@
Fig. (3) shows the $L_{2}$-norm of the Hamiltonian and momentum constraints as a function of grid resolution (focusing on the final iteration at each resolution) for the two systems. In figs. (3(a),3(b)), the convergence test emphasizes on the NS. We observe the constraint violations decrease exponentially for smooth matter field, i.e., K96 EOS. For the SFHo EOS, initially for low resolutions the constraints converge exponentially but for higher resolutions the constraints level off. This behavior is expected; when the grid resolution is coarse the true discontinuities of EOS variables are not seen by the spectral expansion. For fine resolutions all features of EOS variables emerge and while spectral convergence try resolve these features but it is not successful. Hence we see constraints are soon level off and do not decrease as we increase the resolution. We note that, for the tabular EOS (SFHo), the convergence is limited due to the sharp features of $\frac{d\rho_{0}}{dh}$ and $\frac{de}{dh}$ near the surface of the NS (see fig. (2)). Indeed, the overall magnitude of the constraint violations is higher and the rate of convergence is not exponential with the tabular EOS (in contrast to the simple polytrope). Nonetheless, the profile of the convergence tests still suggest convergence up to a limit at high resolutions.
@@ 4.2 Post-Newtonian test
We calculate the binding energy $E_{b}$ of a symmetric BNS system with NS baryon mass $1.4$ and tabulated SLy4 EOS for varied separations of NS. Here $E_{b}=M_{\mathrm{ADM}}-M_{\infty}$ and $M_{\infty}=M^{1}_{\mathrm{TOV}}+M^{2}_{\mathrm{TOV}}$; we use eq. 48 to calculate $M_{\mathrm{ADM}}$ and $M^{1/2}_{\mathrm{TOV}}$ are the corresponding gravitational mass of NS in isolation – found by a TOV solver.
To determine $\Omega^{z}_{\mathrm{BNS}}$ for a specific separation, we employ an iterative approach based on the force balance method described in equation eq. 52. To validate the accuracy of the generated ID against the expected analytical values for significant separations, particularly when the BNS system exhibits quasi-circular motion, we calculate the binding energy $E_{b}$ corresponding to the given $\Omega^{z}_{\mathrm{BNS}}$ utilizing the post-Newtonian formula outlined in [91].
Figure 4: Analytic post-Newtonian curve versus NR curve. The values of $E_{b}$ for given $\Omega^{z}_{\mathrm{BNS}}$ are compared for post-Newtonian method and NR method. For binaries with a large separation post-Newtonian and NR values are matching. For BNS of a close separation there is a slight deviation from post-Newtonian prediction as the system is not fully in quasi-circular status. Here, $\mu={M^{1}_{\mathrm{TOV}}M^{2}_{\mathrm{TOV}}}/{M_{\infty}}$ and the binary has a symmetric baryon mass of $1.4$ for each NS and uses tabulate SLy4 EOS.
Figure (4) illustrates the binding energy of the BNS system across different separations compared to the post-Newtonian data points. In cases of binaries with significant separations, meeting the quasi-circular motion criteria, we observe that the post-Newtonian and NR values align. However, for smaller separations, the NR values of $E_{b}$ are marginally higher.

### v2

@@ 1 Introduction
Simulation of compact binary system in NR often involves two steps. The first step is to find the solution of Einstein-Euler PDE on a hypersurface of the spacetime manifold, namely, constructing constraint satisfying and self-consistent Initial Data (ID) that present the binary system of interest at some time. ID codes such as, COCAL [22, 23, 24], Elliptica [25], FUKA [26], LORENE [27, 28, 29, 30, 31], NRPyElliptic [32], SGRID [33, 34, 35], SpECTRE’s elliptic solver [36, 37], Spells [38, 39, 40, 41], TwoPunctures [42, 43], are developed and utilized for this step.
The next step is to use the ID and simulate the system’s evolution over time; hence we can find the solution over the spacetime of interest. Dynamical evolution codes, for instance, AthenaK [44, 45], BAM [46, 47, 48], BAMPS [49, 50, 51], Dendro-GR [52], Einstein Toolkit [53, 54], ExaHyPE [55] GR-Athena++ [56, 57, 58] GRaM-X [59, 60], GRChombo [61, 62], Nmesh [63], NRPy+ [64], SACRA-MPI [65], Simflowny [66], SpEC [67, 68], SpECTRE [69, 70], and SPHINCS_BSSN [71], among others, are made and employed for this step.
@@
The remainder of paper is organized as follows. In Sec. (2) we explain the mathematical background and formalism we use for ID of BNS systems, in particular, Einstein-Euler equations and EOS in Elliptica. Sec. (3) details the underling algorithms and numerical techniques for construction of BNS’s ID. We present the implementation of tabulated EOS by spline interpolant means and diagnostics in Elliptica. Additionally, we explain the iterative procedure for finding the physical and constraint satisfying ID for BNS systems. In Sec. (4), we present various convergence tests and comparison against post Newtonian answers to showcase the proof of concept for the new version the code. In Sec. (5), we discuss the possible improvements and future work.
Throughout the article, we use geometric units where $G=1$ is the constant of gravity, $c=1$ is the speed of light, and solar mass $M_{\odot}=1$. Additionally, following [77], we absorb the speed of light $c$ in the definition of pressure and total energy density.
@@ 2.1 Einstein-Euler equations
Here $\psi$ is the conformal factor, $\bar{\gamma}_{ij}$ the conformal 3-metric, $A^{ij}$ the traceless part of $K^{ij}$. Moreover, since we are using the Extended Conformal Thin Sandwich (XCTS) formalism [75, 76]
$A^{ij}$ is decomposed as
@@
here, $\Omega^{j}_{\mathrm{NS}}$ is a free parameter to set the spin level, see Sec. (4.3), and $x^{k}_{c}$ denotes the NS coordinate. The irrotational part is shown by the potential $\phi(x,y,z)$ and obeys the following equations. [35]
@@ 3.1 Coordinate system
Figure 1: Grid patches for a BNS system. Left: shown is the $x-y$ plane of the computational grid that is covered by different patches. To fully represent the grid, the other boundary is set to a shorter length. For a production run, the outer boundary is set generally at a radius of $10^{5}\mathrm{\,M_{\odot}}$ from to the center of grid. Right: a zoomed-in view of the $x-y$ plane focusing on the NS regions. A Cartesian box is used around each NS center to avoid coordinate singularities. By using various $\sigma(X,Y)$ values, cubed spherical patches can adapt to different shapes, effectively capturing the shape of the NS surface. This allows for the treatment of matter and vacuum in separate patches.
Fig. (1) illustrates the patches covering the $x-y$ plane of the computational grid used in this work. It is worth noting that the sizes used for the grid and NS are reduced to illustrate the complete computational grid. In practice, we set the outer boundary at a radius of $10^{5}\mathrm{\,M_{\odot}}$.
@@ 3.2 Elliptic solver
Given an elliptic PDE, we linearize the equation to use the Newton-Raphson method (see [25] for a complete description). Thus, the original elliptic PDE becomes a matrix equation, like $Ax=B$, to be solved. Now the challenge is to solve this matrix equation efficiently. To this aim, we use the Schur domain decomposition method [25, 82]. At the core of this method there is a divide and conquer strategy; in this strategy, first the coupled equations are solved, and then the system decomposes into independent (uncoupled) subsystems, amenable for parallel solving. The Schur domain decomposition allows solving the full system using shared memory multiprocessing in which each CPU core is assigned to solve a matrix equations at each patch independently.
The Schur domain decomposition method arranges the Jacobian matrix in Newton-Raphson method into two parts. The first part is associated with all uncoupled equations and the second part comprises the coupled equations. These couplings between equations occur due to the inter-BC at the interfaces of adjacent patches [25]. As such, the system of matrix equations can be seen as two equations with two unknowns as follows
@@
Therefore, we first solve for $\tilde{w}$, and then we use it to solve for $\tilde{v}$. Accordingly, in Elliptica, we solve Eq. (35) like:
@@
As we can see eq (36) only involves the unknown $\tilde{w}$, and it can be summarized as
@@ 3.4 Tabulated equations of state
A number of slight modifications may be made to the EOS in preprocessing in order to make it more amenable to interpolation. These modifications are necessary for both the physical consistency of the EOS and to facilitate the convergence of the solution.
The first modification involves adjusting specific enthalpy at the NS surface for the EOS tables in which the specific enthalpy falls below $1$. To this aim, at the minimum pressure, which we consider occurring at the beginning of the table, we scale all $\rho_{0}$ points by the multiplicative constant $\eta=h_{\mathrm{min}}$ corresponding to the lowest value of $h$ in the table. Accordingly, $h$ points are scaled by the value $\eta^{-1}$, i.e:
@@
This alteration can be carried through to dynamical evolution codes by e.g. using the same EOS table or by scaling the rest-mass density, at the minimum pressure, by the same factor used in the preprocessing step.
In addition, several other features of tabular EOS may pose problems for both interpolation and the use of specific enthalpy as the independent thermodynamic variable. Among these problems is the presence of a region near the surface of the NS where both $\frac{de}{dh}$ and $\frac{d\rho_{0}}{dh}$ diverge. This region is common to many EOS with a ‘crust’, including both tabular EOS such as SFHo [86] and piecewise polytropics [87]. While this region is unavoidable if we wish to accurately represent the equation of state via specific enthalpy, cf. [88, 89], it also limits the accuracy of the solution within the NS (see Fig. (3)).
@@
Figure 2: Comparison of changes with respect to the specific enthalpy ($h$) in the rest mass density ($\rho_{0}$) near the NS surface for three EOS: K96 (single polytrope), SLy (piecewise polytrope), and SFHo (table). A lower specific enthalpy corresponds to a point closer to the NS surface, with $h=1$ defining the surface itself. $\rho_{0}$ is in geometric unit and $h$ is dimensionless.
Fig. (2) shows the resultant tabulated SFHo EOS after re-griding using a low-degree Hermite interpolant with comparison to EOS of K96 and SLy. The K96 EOS is a single polytrope, corresponding to $K=96.7$ and $\Gamma=2$ [90]. SLy is an approximate piecewise polytrope of tabulated SLy4, and therefore quantities such as $\frac{d\rho_{0}}{dh}$ are not continuous in certain regions corresponding to transitions between piecewise segments [91]. SFHo, here restricted to the zero-temperature regime, shows similar discontinuities [86].
Examining the functions relevant to the ID generation, e.g. $\rho_{0}$ and $e$, we see that the single-polytrope model is considerably simpler and more amenable to numerical solution than the models with discontinuities. The difference is typically relevant to the surface of the NS (where $h=1$), as shown in the nearly vertical profile of $\frac{d\rho_{0}}{dh}$ in Fig. (2). The sharp changes in $\frac{d\rho_{0}}{dh}$ with the piecewise polytropic and tabular EOS are not perfectly captured by the grid, and thus they limit the convergence of the constraint violations (see Sec. (4.1)).
There are several potential ways to improve the convergence properties of the code in this area. For example, one could apply filtering to the spectral coefficients [92] of $\rho_{0}$ or implement a routine to identify the location of discontinuities, thereby creating patches where the discontinuity occurs at the boundary between those patches. Additionally, using spectral EOS [88, 89, 93], and imposing continuity at the piecewise polytrope EOS [87] have been shown to enhance the continuity of the solutions. The implementation and comparison of these methods are left for future work.
@@ 3.5 Initial data construction
Step 1: We solve the elliptic equations iteratively in this specific order: first Eq. (25) for the matter field, and then Eqs. (17), (18), and (19) for the metric fields. This approach has been found to enhance the solution’s stability, particularly when solving the matter field first. The order of solving the metric fields themselves, however, appears to be less critical. Moreover, during each iteration step, we focus on solving a single elliptic equation while keeping the other fields fixed. In essence, the fixed fields act as source terms influencing the equation being solved. Finally, within the Newton-Raphson iterative method, we perform only one update step per field, and then incorporating the newly solved value into the source terms for the next equation. In the following $\Xi$ denotes a field from the set $\{\phi,\psi,\alpha\psi,B^{i}\}$.
@@
Step 3: After updating all $\Xi$ fields, we see the baryon mass of each NS deviates, often by a few percents, from the target value. Additionally, for spinning NS, since at Step 0 we used TOV solution, and then we added the spin vector to the NS, it is not surprising to observe the baryon mass is different from the prescribed target value. Moreover, the starting resolution is often coarse, and again the baryon mass may change as we go to a higher resolution. If we do not account for these changes, the baryon mass deviates even further at later iterations and may lead to a code crash. To adjust the baryon mass, we note that $\rho_{0}=\rho_{0}(h)$, and as shown in Eq. (31) the specific enthalpy depends on a constant $C$, i.e., $h=h(C)$, which implies $\rho_{0}=\rho_{0}(C)$. Therefore by using a root finder, we find the value of $C$ in Eq. (40) such that the baryon mass remains the same as the target value.
@@
Step 6: Finding the orbital angular velocity, $\Omega^{z}_{\mathrm{BNS}}$, for a quasi-circular orbit demands full solution of general relativity. Since we start with a Newtonian approximation for this value we need to refine it. Following [94], so called the force balance method, we first compute $\partial_{i}\ln\Gamma$ at each NS center, where $\Gamma$ computed as
@@
Step 7: At this point the specific enthalpy profile and hence NS surface are changed. We need to find the new location of NS surface for adjusting cubed spherical patches that are covering the NS – so we can separate matter and vacuum into different patches. Since some parts of the NS may need to extend to the patches that are currently covering vacuum, we extrapolate $h$ into these patches, so the root finder can find where $h=1$. To this aim we use the following formula to extrapolate specific enthalpy to the vacuum
@@
here, $r$ denotes the coordinate distance from the NS center, and $c_{0}$ is a constant (typically $0.01$). We find the values of $a$, and $b$ by demanding $C^{1}$ continuity across the NS surface. Additionally, we use the same Eq. (53) to extrapolate $\phi$ field outside the NS. This step is required when the NS surface is expanded, and we want to interpolate $\phi$ from the current grid to a new grid.
@@
Step 11: We monitor the Hamiltonian and momentum constraints, Eqs. (55) and (56). We restart from Step 1 unless the constraints reach their truncation error and are level off. In this case, when the constraints reach a plateau, we stop the iterative process if there is no higher resolution demanded; otherwise we go to the next resolution and start from Step 1.
@@ 4.1 Convergence test
For the convergence test, we generate ID for two symmetric BNS systems: one with polytropic K96 EOS and the other with tabulated SFHo EOS. In these systems, the NS have baryon mass $1.4$ with no spin, and their separation is $50M_{\odot}$.
@@
Fig. (3) shows the $L_{2}$-norm of the Hamiltonian and momentum constraints as a function of grid resolution (focusing on the final iteration at each resolution) for the two systems. In figs. (3(a),3(b)), the convergence test emphasizes on the NS. We observe the constraint violations decrease exponentially for smooth matter field, i.e., K96 EOS. For the SFHo EOS, initially for low resolutions the constraints converge exponentially but for higher resolutions the constraints level off. This behavior is expected; when the grid resolution is coarse the true discontinuities of EOS variables are not seen by the spectral expansion. For fine resolutions all features of EOS variables emerge and while spectral convergence try to resolve these features, but it is not successful. Hence, we see constraints are soon level off and do not decrease as we increase the resolution. We note that, for the tabular EOS (SFHo), the convergence is limited due to the sharp features of $\frac{d\rho_{0}}{dh}$ near the surface of the NS (see Fig. (2)). Indeed, the overall magnitude of the constraint violations is higher and the rate of convergence is not exponential with the tabular EOS (in contrast to the simple polytrope). Nonetheless, the profile of the convergence tests still suggest convergence up to a limit at high resolutions.
@@ 4.2 Post-Newtonian test
We calculate the binding energy $E_{b}$ of a symmetric non-spinning BNS system with NS baryon mass $1.4$ and tabulated SLy4 EOS for varied separations of NS. Here $E_{b}=M_{\mathrm{ADM}}-M_{\infty}$ and $M_{\infty}=M^{1}_{\mathrm{TOV}}+M^{2}_{\mathrm{TOV}}$; we use Eq. (48) to calculate $M_{\mathrm{ADM}}$ and $M^{1/2}_{\mathrm{TOV}}$ are the corresponding gravitational mass of NS in isolation – found by a TOV solver.
To determine $\Omega^{z}_{\mathrm{BNS}}$ for a specific separation, we employ an iterative approach based on the force balance method described in Eq. (52). To validate the accuracy of the generated ID against the expected analytical values for significant separations, particularly when the BNS system exhibits quasi-circular motion, we calculate the binding energy $E_{b}$ corresponding to the given $\Omega^{z}_{\mathrm{BNS}}$ utilizing a post-Newtonian formula, the 4 post-Newtonian order approximation, outlined in [95], Sec. 7.4. Specifically, to compute the post-Newtonian values of $E_{b}$, we need the reduced mass $\mu=M^{1}_{\mathrm{TOV}}M^{2}_{\mathrm{TOV}}/{M_{\infty}}$, the symmetric mass $\nu=\mu/M_{\infty}$, and the dimensionless orbital velocity $x^{3/2}=\Omega^{z}_{\mathrm{BNS}}M_{\infty}$.
Figure 4: Analytic post-Newtonian (PN) curve versus NR curve. The values of $E_{b}$ for given $\Omega^{z}_{\mathrm{BNS}}$ are compared for post-Newtonian method and NR method. For binaries with a large separation post-Newtonian and NR values are matching. For BNS of a close separation there is a slight deviation from post-Newtonian prediction as the system is not fully in quasi-circular status. Here, $\mu={M^{1}_{\mathrm{TOV}}M^{2}_{\mathrm{TOV}}}/{M_{\infty}}$ and the binary has a symmetric baryon mass of $1.4$ for each NS and uses tabulated SLy4 EOS.
Fig. (4) illustrates the binding energy of the BNS system across different separations compared to the post-Newtonian data points. In cases of binaries with significant separations, meeting the quasi-circular motion criteria, we observe that the post-Newtonian and NR values align. However, for smaller separations, the NR values of $E_{b}$ are marginally higher.